# DINO-WM: World Models on Pre-trained Visual Features enable Zero-shot Planning

### v1

@@ A.1 Environments and Dataset Generation
    Point Maze: In this environment entroduced by Fu et al. (2021), the task is for a force-actuated 2-DoF ball in the catesian directions x and y to reach a target goal. The agent’s dynamics incorporate physical properties such as velocity, acceleration, and inertia, making the movement realistic. We customize the environment by altering the maze configuration to test the model’s generalization ability in unseen situations. We generate 2000 fully random trajectories to train our world models.
@@
    Push-T: This environment introduced by Chi et al. (2024) features a pusher agent interacting with a T-shaped block. The goal is to guide both the agent and the T-block from a randomly initialized state to a known feasible target configuration within 25 steps. The task requires both the agent and the T to match the target locations. Unlike previous setups, the fixed green T no longer represents the target position for the T-block but serves purely as a visual anchor for reference. Success requires precise understanding of the contact-rich dynamics between the agent and the object, making it a challenging test for visuomotor control and object manipulation. Additionally, we introduce variations by altering the shape and color of the object to assess the model’s capability to adapt to novel tasks. We generate a dataset of 18500 samples replayed the original released expert trajectories with various level of noise and evaluate the model’s performance across all different shapes to assess its adaptability.
@@
    Wall: This custom 2D navigation environment features two rooms separated by a wall with a door. The agent’s task is to navigate from a randomized starting location in one room to a goal in the other, passing through the door. We present a variant where wall and door positions are randomized, testing the model’s generalization to novel configurations. For the fixed wall setting, we train on a fully random dataset of 2000 trajectories each with 50 time steps. For the variant with multiple training environment configurations, we generate 10240 random trajectories.
@@
    Rope Manipulation: Introduced in Zhang et al. (2024), this task is simulated with Nvidia Flex Zhang et al. (2024) and consists of an XArm interacting with a soft rope placed on a tabletop. The objective is to move the rope from an arbitrary starting configuration to a goal configuration specified at inference time. For training, we generate a random dataset of 1000 trajectories of 20 time steps of random actions from random starting positions, while testing involves goal configurations set from varied initial positions, incorporating random variations in orientation and spatial displacement.
@@ A.3 Pretraining features
    R3M: A ResNet-18 model pre-trained on a wide range of real-world human manipulation videos Nair et al. (2022).
@@
    ImageNet: A ResNet-18 model pre-trained on the ImageNet-1K dataset Russakovsky et al. (2015).
@@
A.4 Planning Optimization
@@
A.4.1 Model Predictive Control with Cross-Entropy Method
@@ A.4.2 Gradient Descent:
A.4.3 Planning Results
Here we present the full planning performance using various planning optimization methods. CEM denotes the setting where we use CEM to optimize a sequence of actions, and execute those actions in the environment without any correction or replan. Similarly, GD denotes optimizing with gradient decent and execute all planned actions at once in an open-loop way. MPC denotes allowing replan and receding horizon with CEM for optimization.
@@ A.4.3 Planning Results
A.5 Comparison with Action-Conditioned Generative Models
We compare DINO-WM with a variant of AVDC, where the diffusion model is trained to generate the next observation $o_{t+1}$ conditioned on the current observation $o_{t}$ and action $a_{t}$, rather than generating an entire sequence of observations at once conditioned on a text goal. We then present openloop rollout results on validation trajectories using this action-conditioned diffusion model, with visualizations shown in Figure 8. It can be seen that the action-conditioned diffusion model diverges from the ground truth observations over long-term predictions, making it insufficient for accurate task planning.
Figure 8: Openloop rollout on PushT with DINO-WM and action-conditioned AVDC (AVDC-AC). For each trajectory, the model is given the first frame as well as sequence of actions. The world models performs openloop rollout with these actions.
A.6 Hyperparameters and implementation
@@
Table 6: Environment-dependent hyperparameters for DINO-WM training
$H$
PointMaze 3
Push-T 3
Wall 1
Rope 1
Granular 1
@@
A.7 Additional Planning Visualizations
We present additional visualizations for planning with DINO-WM. In this setting, all planning instances share the same initial observations but have different goal observations to demonstrate DINO-WM’s generalization capabilities in planning. We show trajectory pairs to compare the environment’s observations after executing a sequence of planned actions with DINO-WM’s imagined trajectories. The left-most column denotes the initial observations, and the right-most shaded column denotes the goal observations. Each pair of rows represents a planning instance: the top (shaded) row shows the environment’s observation after executing 25 planned actions, and the bottom row shows the world model’s imagined observations.

### v2

@@ A.1 Environments and Dataset Generation
    PointMaze: In this environment introduced by (Fu et al., 2021), the task is for a force-actuated 2-DoF ball in the Cartesian directions $x$ and $y$ to reach a target goal. The agent’s dynamics incorporate physical properties such as velocity, acceleration, and inertia, making the movement realistic. We generate 2000 fully random trajectories to train our world models. We refer to this task as Maze for brevity in our tables.
@@
    Wall: This custom 2D navigation environment features two rooms separated by a wall with a door. The agent’s task is to navigate from a randomized starting location in one room to a goal in the other, passing through the door. We present a variant where wall and door positions are randomized, testing the model’s generalization to novel configurations. For the fixed wall setting, we train on a fully random dataset of 1920 trajectories each with 50 time steps. For the variant with multiple training environment configurations, we generate 10240 random trajectories.
@@
    Reacher: A continuous control task from DeepMind Control Suite (Tassa et al., 2018), where a 2-joint robotic arm reaches a target in 2D space. We increase difficulty by requiring the entire arm, not just the end-effector, to match arbitrary target poses. To train the world model, we generate 3000 trajectories with 100 steps. We refer to this task as Reach for brevity in our tables.
@@
    Push-T: This environment introduced by (Chi et al., 2024) features a pusher agent interacting with a T-shaped block. The goal is to guide both the agent and the T-block from a randomly initialized state to a known feasible target configuration within 25 steps. The task requires both the agent and the T to match the target locations. Unlike previous setups, the fixed green T no longer represents the target position for the T-block but serves purely as a visual anchor for reference. Success requires precise understanding of the contact-rich dynamics between the agent and the object, making it a challenging test for visuomotor control and object manipulation. We generate a dataset of 18500 samples replayed the original released expert trajectories with various level of noise. Additionally, we introduce variations by altering the shape and color of the object to assess the model’s capability to adapt to novel tasks. For this variant, we generate 20000 randomly sampled trajectories with 100 steps.
@@
    Rope Manipulation: Introduced in (Zhang et al., 2024), this task is simulated with Nvidia Flex (Zhang et al., 2024) and consists of an XArm interacting with a soft rope placed on a tabletop. The objective is to move the rope from an arbitrary starting configuration to a goal configuration specified at test time. For training, we generate a random dataset of 1000 trajectories of 20 time steps of random actions from random starting positions, while testing involves goal configurations set from varied initial positions, incorporating random variations in orientation and spatial displacement.
  6. f)
@@ A.3 Pretraining features
    R3M: A ResNet-18 model pre-trained on a wide range of real-world human manipulation videos (Nair et al., 2022).
@@
    ImageNet: A ResNet-18 model pre-trained on the ImageNet-1K dataset (Russakovsky et al., 2015).
@@
A.4 Ablations
A.4.1 Scaling Laws of DINO-WM
To analyze the scaling behavior of DINO-WM, we trained world models and performed planning using datasets of varying sizes, ranging from 200 to 18500 trajectories on the PushT environment. Our results demonstrate a clear trend: as the dataset size increases, both the quality of the world model’s predictions and the performance of the planned behavior improve significantly. Larger datasets enable the world model to capture more diverse dynamics and nuances of the environment, leading to more accurate predictions and better-informed planning.
Table 5: Planning performance and prediction quality on PushT with DINO-WM trained on datasets of various sizes. SSIM and LPIPS are measured on the predicted future latents after decoding. We observe consistent improvement in performance as we increase the dataset size.
Dataset Size SR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ n=200 0.08 0.949 0.056 n=1000 0.48 0.973 0.013 n=5000 0.72 0.981 0.007 n=10000 0.88 0.984 0.006 n=18500 0.92 0.987 0.005
A.4.2 DINO-WM with vs. without Causal Attention Mask
We introduce a causal attention mask in Section 3.1.2. We ablate this choice on PushT by training DINO-WM with and without this causal attention mask with varying history length $h$, such that the model takes in input $o_{t-h+1},o_{t-h+2},...o_{t}$, and output $o_{t-h+2},...o_{t+1}$. For models with mask, the model can only attend to past observations for predicting each $o_{t}$, whereas in the w/o mask case, predicting any observation in the output sequence can attend to the entire input sequence of observations. We show planning success rate on our PushT settings in Table 6. When $h=1$ where the model with and without this causal mask is equivalent, both models get decent and equivalent success rate. However, as we increase the history length, we see a rapid drop in the w/o mask case, since the model can cheat during training by attending to future frames, which are not available at test time. Adding the causal mask solves this issue, and we observe improvement in performance as longer history could better capture dynamics information like velocity, acceleration, and object momentum.
Table 6: Comparison of DINO-WM with and without causal attention mask on PushT. We train models with varying history $h$, representing the number of past observations the model takes as input.
$h=1$ $h=2$ $h=3$ w/o mask 0.76 0.36 0.08 with mask 0.76 0.88 0.92
A.4.3 DINO-WM with Reconstruction Loss
While DINO-WM eliminates the need to train world models with a pixel reconstruction loss—avoiding the risk of learning features irrelevant to downstream tasks—we conduct an ablation study where the predictor is trained using a reconstruction loss propagated from the decoder. As shown in Table 7, this approach performs reasonably well on the PushT task but falls slightly short of our method, where the predictor is trained entirely independently of the decoder. This underscores the advantage of disentangling feature learning from reconstruction objectives.
Table 7: Comparison of DINO-WM trained with and without loss from the decoder on PushT, highlighting the advantage of disentangling feature learning from reconstruction objectives.
Success Rate w/o decoder loss 0.92 with decoder loss 0.80
A.5 Planning Optimization
@@
A.5.1 Model Predictive Control with Cross-Entropy Method
@@ A.5.2 Gradient Descent:
A.5.3 Planning Results
Here we present the full planning performance using various planning optimization methods. CEM denotes the setting where we use CEM to optimize a sequence of actions, and execute those actions in the environment without any correction or replan. Similarly, GD denotes optimizing with gradient decent and executing all planned actions at once in an open-loop way. MPC denotes allowing replan and receding horizon with CEM for optimization.
@@ A.5.3 Planning Results
A.6 Comparison with Action-Conditioned Generative Models
We compare DINO-WM with a variant of AVDC, where the diffusion model is trained to generate the next observation $o_{t+1}$ conditioned on the current observation $o_{t}$ and action $a_{t}$, rather than generating an entire sequence of observations at once conditioned on a text goal. We then present open-loop rollout results on validation trajectories using this action-conditioned diffusion model, with visualizations shown in Figure 7. It can be seen that the action-conditioned diffusion model diverges from the ground truth observations over long-term predictions, making it insufficient for accurate task planning.
Figure 7: Open-loop rollout on PushT with DINO-WM and action-conditioned AVDC (AVDC-AC). For each trajectory, the model is given the first frame as well as sequence of actions. The world models performs open-loop rollout with these actions.
A.7 Decoding the Latents: LPIPS and SSIM Metrics
We report two key metrics: Structural Similarity Index (SSIM) (Wang et al., 2004) and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) on the reconstruction of world models’ predicted future states across four of our more challenging environments. SSIM measures the perceived quality of images by evaluating structural information and luminance consistency between predicted and ground-truth images, with higher values indicating greater similarity. LPIPS assesses perceptual similarity by comparing deep representations of images, with lower scores reflecting closer visual similarity.
LPIPS $\downarrow$ SSIM $\uparrow$
Method PushT Wall Rope Granular PushT Wall Rope Granular
R3M 0.045 0.008 0.023 0.080 0.956 0.994 0.982 0.917
ResNet 0.063 0.002 0.025 0.080 0.950 0.996 0.980 0.915
DINO CLS 0.039 0.004 0.029 0.086 0.973 0.996 0.980 0.912
AVDC 0.046 0.030 0.060 0.106 0.959 0.983 0.979 0.909
Ours 0.007 0.0016 0.009 0.035 0.985 0.997 0.985 0.940
A.8 Inference Time
Inference time is a critical factor when deploying a model for real-time decision-making. Table 10 reports the time required on an NVIDIA A6000 GPU for a single inference step, the environment rollout time for advancing one step in the simulator, and the overall planning time for generating an optimal action sequence using the Cross-Entropy Method (CEM). The inference time of DINO-WM remains constant across environments due to the fixed model size and input image resolution, resulting in significant speedup over traditional simulation rollouts. Notably, in environments with high computational demands, such as deformable object manipulation, simulation rollouts require several seconds per step while DINO-WM enables rapid inference and efficient planning. Planning time is measured with CEM using 100 samples per iteration and 10 optimization steps, demonstrating that DINO-WM can achieve feasible planning times while maintaining accuracy and adaptability across tasks.
Table 10: Inference time and planning time for DINO-WM. Inference time represents the time required for a single forward pass for one step, while environment rollout time measures the simulator’s speed for advancing one step. Planning time corresponds to Cross-Entropy Method (CEM) with 100 samples per iteration and 10 optimization steps.
Metric Time (s)
Inference (Batch 32) 0.014
Simulation Rollout (Batch 1) 3.0
Planning (CEM, 100x10) 53.0
A.9 Hyperparameters and implementation
@@
The world model architecture is consistent across all environments. We use an encoder based on DINOv2, which extracts features with a shape of $(14\times 14,384)$ from input images resized to $196\times 196$ pixels. The ViT backbone has a depth of 6, 16 attention heads, and an MLP dimension of 2048, amounting to approximately 19M parameters.
To ensure the prediction task is meaningful, as nearby observations can be highly similar, we introduce a frameskip parameter during data processing. This parameter specifies how far into the future the model is predicting. The frameskip values for each environment are provided in Table 12.
Table 11: Environment-dependent hyperparameters for DINO-WM training. We report the number of trajectories in the dataset under Dataset Size, and the length of trajectories under Traj. Len.
$H$ Frameskip Dataset Size Traj. Len.
PointMaze 3 5 2000 100
Reacher 3 5 3000 100
Push-T 3 5 18500 100-300
PushObj 3 5 20000 100
Wall 1 5 1920 50
WallRandom 1 5 10240 50
Rope 1 1 1000 5
Granular 1 1 1000 5
@@
A.10 Additional Planning Visualizations
We show visualizations of planning instances for DINO-WM and our baselines in Figure 8. For comparison, we show the best performing world models DINO CLS and DreamerV3. We also show visualizations of DINO-WM on all tasks in Figure 9. For each environment, the top (shaded) row shows the environment’s observation after executing the planned actions, and the bottom row shows the world model’s imagined observations.
Figure 9: Trajectories planned with DINO-WM on all six environments. For each environment, the top (shaded) row shows the environment’s observation after executing the planned actions, and the bottom row shows the world model’s imagined observations.
To demonstrate DINO-WM’s ability to generalize to different goals at test time, we show additional visualizations for DINO-WM when provided with the same initial observation but different goal observations in Figure 10 and Figure 11. Similarly, we show trajectory pairs to compare the environment’s observations (top shaded rows) after executing a sequence of planned actions with DINO-WM’s imagined trajectories (bottom rows). The left-most column denotes the initial observations, and the right-most shaded column denotes the goal observations.